### v1

# Mitigating the Linguistic Gap with Phonemic Representations for Robust Multilingual Language Understanding

Abstract
Approaches to improving multilingual language understanding often require multiple languages during the training phase, rely on complicated training techniques, and—importantly—struggle with significant performance gaps between high-resource and low-resource languages. We hypothesize that the performance gaps between languages are affected by linguistic gaps between those languages and provide a novel solution for robust multilingual language modeling by employing phonemic representations (specifically, using phonemes as input tokens to LMs rather than subwords). We present quantitative evidence from three cross-lingual tasks that demonstrate the effectiveness of phonemic representation, which is further justified by a theoretical analysis of the cross-lingual performance gap [^1] Our code is available here: https://anonymous.4open.science/r/ipa-for-multilingual-nlu/README.md.
Mitigating the Linguistic Gap with Phonemic Representations
for Robust Multilingual Language Understanding
Hae Ji Jung Korea University gpwl0709@korea.ac.kr                        Changdae Oh University of Seoul                        Jooeon Kang Sogang University                        Jimin Sohn GIST
Kyungwoo Song Yonsei University                        Jinkyu Kim Korea University                        David R. Mortensen Carnegie Mellon University dmortens@cs.cmu.edu
1 Introduction
In an era of large language models, natural language processing has promised to bring us together. However, there are gaps between how human language technologies perform in some languages (like English) and others (most of the languages of the world). Language technologies exhibit performance gaps between languages. To some degree, this is due to differences in resourcedness, but we hypothesize that linguistic gaps—the chasms that separate languages from one another structurally and lexically—also play a role. We further hypothesize that techniques that reduce linguistic gaps will also reduce performance gaps. In this paper, we focus on one such technique: representing languages phonetically using the International Phonetic Alphabet (IPA).
Figure 1 illustrates our motivation for using IPA as a universal language representation. The key contributions of this paper are as follows:
  • We demonstrate the advantage of using phonemic representation (i.e., IPA) for language modeling, particularly as a robust multilingual representation.
  • We empirically validate the effectiveness of phonemic representations by comparing the performance gap and linguistic gap across languages with subword or character representations.
  • We further explain the empirical observations with theoretical analysis from domain generalization literature, by treating the linguistic gap as the domain gap caused by lexical and syntactic disparities.
Figure 1: Example of word, character, and phoneme units for a sentence (English and Korean).
2 Related Work
Cross-lingual Transfer Learning
approaches aim to transfer knowledge from one language to enhance NLP performance in another. While adversarial training methods Lample et al. (2018); Chen et al. (2018, 2021); Yu et al. (2023) attempt to learn a language-agnostic representation, their training can be unstable Balcan et al. (2023). Non-adversarial joint learning approaches Cotterell and Heigold (2017); Gu et al. (2018); Wei et al. (2021); Zheng et al. (2021) also present promising results, but necessitate significant multilingual data. Recently, Yang et al. (2022) revealed the empirical correlation between cross-lingual transferability and representation discrepancy, yet they do not provide any theoretical statement to justify such numerical analysis, which is presented in this work.
Multilingual Language Models (MLLMs)
have been rapidly advanced with a popularized large-scale pre-training and fine-tuning paradigm Devlin et al. (2018); Conneau and Lample (2019); Conneau et al. (2020); Chi et al. (2021); Workshop et al. (2022); Wei et al. (2023). Although MLLMs have demonstrated remarkable potential in diverse multilingual tasks, there is a significant performance gap between different languages, especially in the case of high-resource language versus low-resource language Wu and Dredze (2020); Zhao et al. (2023). In this work, we propose a novel paradigm for robust multilingual language modeling: IPA as a universal language representation.
Phonemic Language Representation
has been applied in previous studies Bharadwaj et al. (2016); Chaudhary et al. (2018); Dalmia et al. (2019); Hu et al. (2019); Nguyen et al. (2023), but primarily for speech recognition tasks or where they were not focused on the linguistic gap between languages. Our work revisits phonemic representation for MLLMs to mitigate the linguistic gap for the first time.
3 Experimental Setup
In this section, we briefly describe the experiment setup in terms of models, datasets, and tasks. Refer to Appendix A for further details.
3.1 Models and Data Preprocessing
We compare LMs with three different types of language representation: subword, character, and phoneme. All models222pre-trained weights were obtained from https://huggingface.co/models are pre-trained on multilingual data that covers around 100 languages from Wikipedia dump files. We employ an off-the-shelf MLLMs, multilingual BERT (mBERT; Devlin et al. (2018)). For a subword-based model which is trained with byte-pair encoding. For a character-based model, we utilize CANINE Clark et al. (2022), which is a tokenization-free LM that directly maps each character to its codepoint by hashing. It is pre-trained on the same data and training objectives as mBERT. For a phoneme-based model, we adopt XPhoneBERT Nguyen et al. (2023), which has the same model architecture as mBERT.
While character-level models are known to better generalize to low-resource languages Clark et al. (2022), their general performance falls behind subword-based models. For a fair comparison between the representations, we primarily compare phoneme-based model to character-based one instead of directly comparing it to the subword-based model (i.e., mBERT), and leave further improvements of overall performance as future work.
Preprocessing.
In order to prepare inputs for a phoneme-based model, we employed G2P (Grapheme-to-Phoneme) conversion to obtain an IPA version of the input. This conversion was done with Epitran [^3] https://github.com/dmort27/epitran Mortensen et al. (2018), an external tool for G2P conversion. After converting to IPA, we inserted white space between every character to make it compatible with XPhoneBERT’s tokenizer.
3.2 Downstream Tasks
We adopted the cross-lingual generalization benchmark tasks suggested in XTREME Hu et al. (2020). Among them, we selected two types of tasks; classification and structured prediction. For evaluation languages, we choose eight languages of three groups—high-resource (eng, deu, fra), low-resource (urd, hin, swa), and typologically distinct (kor, ukr)—to analyze the impact of phonemic representation with respect to each category.
Sentence-level Classification.
We employ XNLI Conneau et al. (2018) dataset, which is a representative benchmark for the natural language inference task on the cross-lingual generalization setting.
Token-level Classification.
We choose POS tagging and NER as our testbed for structured prediction tasks, both requiring predicting labels for each word in sentences. We utilize the corpora from Universal Dependencies [^4] https://universaldependencies.org/ , v2.13, 148 languages, released Nov 15, 2023. for POS tagging, and WikiAnn Pan et al. (2017) with train, dev, test splits following Rahimi et al. (2019) for NER.
4 Results
Method Language
ENG SWA URD
($\Delta$ from ENG) ($\Delta$ from ENG)
Subword 80.80 62.93 (17.87) 61.57 (19.23)
Character 75.02 59.72 (15.30) 56.55 (18.46)
Phoneme 71.89 60.88 (11.00) 56.10 (15.78)
Table 1: Accuracy (%) on XNLI task. ENG, SWA, URD refer to English, Swahili, and Urdu, respectively. Phonemic representation shows relatively small performance gaps compared to other representations.
Method Language Performance gap Linguistic Gap
EN DE FR KO UK UR HI SW Std. ($\downarrow$) Mean diff. ($\downarrow$) Mean CKA ($\uparrow$)
Named Entity Recognition
Character 88.80 77.99 90.87 80.17 85.61 75.88 73.33 74.84 6.68 0.10 0.34
Phoneme 91.36 81.36 91.93 79.33 83.67 76.41 82.19 78.06 5.79 0.08 0.43
Part-of-Speech Tagging
Character 73.54 81.20 80.55 70.1 84.28 93.15 - - 8.14 0.12 0.30
Phoneme 70.50 78.68 78.10 77.35 83.75 93.96 - - 7.88 0.12 0.52
Table 2: Performance of POS tagging and NER across different languages. Std. refers to the standard deviation of the scores across the languages, and Mean diff. indicates average pairwise difference of F1 scores. Mean CKA represents the average linguistic gap between languages.
Figure 2: Linguistic gaps across languages. Upper and lower triangular elements indicate pairwise linguistic gaps derived with phoneme-based model and character-based model, respectively. Lighter color indicates larger CKA score, which means smaller discrepancy. Upper triangular elements show relatively lighter colors, implying smaller discrepancies across languages.
Figure 3: Qualitative analysis of performance gap (difference of accruacy) on XNLI task. (Left) the absolute difference between performance across two languages, (center) centered kernel alignment (CKA) scores to measure cross-lingual embedding similarity, and (right) Sinkhorn distance on the output probability space. Phonemic representation shows relatively small performance gaps w.r.t. EN $\leftrightarrow$ SW and EN $\leftrightarrow$ UR, and these gaps are correlated with similarity and discrepancy on the embedding space (CKA) and logit space (Sinkhorn distance).
Phoneme-based Model on Low-Resource Languages.
The phoneme-based model shows promising results compared to other models, especially for low-resource languages. As shown in Table 1, the phoneme-based model has the smallest performance gap between English and other low-resource languages – swa, urd. Furthermore, while subword-based mBERT achieves the highest scores, the performance disparity across models narrows when it comes to low-resource languages. Table 2 also suggests that the phoneme-based model exhibits superiority in addressing low-resource languages. For NER, on languages like urd, hin, and swa, the phoneme-based model significantly outperforms character-based model, highlighting the capability of the phoneme-based model’s generalization over the low-resource languages.
Performance Gap Across Languages.
We observe that the phoneme-based model performs the most consistently across languages. The leftmost panel in Figure 3, shows how each language representation results in performance gaps across different languages. Here, the phoneme-based model comes with the lowest performance gap. Table 2 also shows the phoneme-based model’s robustness in terms of performance gap across languages. From the table, standard deviation (Std.) and mean difference (Mean diff.) both indicate how the results for all languages differ from each other. The phoneme-based model ends up with a smaller standard deviation in both NER and POS tasks.
Linguistic Gap of Different Representations.
To analyze the potential of phonemes as a robust representation for multilingual language modeling, we check the gap between different languages over different representations. As in Yang et al. (2022), we use centered kernel alignment (CKA) to compute the representation similarity. Figure 3 shows that phonemic representation shows higher CKA and lower Sinkhorn distance compared to other representations, meaning that the phonemic representations from different languages are relatively closer to each other than those of subword or character representations.
Figure 2 also illustrates the linguistic gap between languages with their pairwise similarity. After fine-tuning each model on downstream tasks, we compute the discrepancy between different languages using held-out parallel data. It can be clearly stated that the upper triangle part (which corresponds to phonemic representations) of the map has large values that imply a smaller linguistic gap, contributing to a smaller performance gap.
Theoretical Analysis.
We aim to diminish the performance gap between different languages by adopting IPA as a universal language representation. Motivated by domain adaptation literature Kifer et al. (2004); Ben-David et al. (2010), we present a theoretical justification of IPA for robust multilingual modeling by deriving a bound for cross-lingual performance gap.
Let $\mathcal{D}$ denote a domain as a distribution over text feature input $\mathcal{X}$, such as the sequence of word embeddings or one-hot vectors, and a labeling function $f:\mathcal{X}\rightarrow\{0,1\}$. Assuming a binary classification task, our goal is to learn a hypothesis $h:\mathcal{X}\rightarrow\{0,1\}$ that is expected to minimize a risk $\varepsilon_{D}(h,f):=\mathbb{E}_{x\sim\mathcal{D}}[\mathbb{I}(f(x)\neq h(x))]$ and has a small risk-deviation over two domains $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$. Then, to formalize the cross-lingual performance gap, we first need a discrepancy measure between two languages. By following Ben-David et al. (2010), we adopt $\mathcal{H}$-divergence (See Appendix B for its definition) to quantify the distance between two language distributions.
Now, based on Lemma 1 and 3 of Ben-David et al. (2010), we make reasoning on performance gap over different language domains.
Theorem 4.1.
Let $h:\mathcal{X}\rightarrow[0,1]$ be a real-valued function in a hypothesis class $\mathcal{H}$ with a pseudo dimension $\mathcal{P}dim(\mathcal{H})=d$. If $\hat{\mathcal{D}}_{A}$ and $\hat{\mathcal{D}}_{B}$ are the empirical distribution constructed by $n$-size i.i.d. samples, drawn from $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ respectively, then for any $\delta\in(0,1)$, and for all $h$, the bound below hold with probability at least $1-\delta$.
$|\varepsilon_{\mathcal{D}_{A}}(h,f)-\varepsilon_{\mathcal{D}_{B}}(h,f)|\leq{1\over 2}d_{\mathcal{H}\Delta\mathcal{H}}(\hat{\mathcal{D}}_{A},\hat{\mathcal{D}}_{B})$
$+2\sqrt{{{d\log(2n)+\log(2/\delta)}\over n}}$
where $\mathcal{H}\Delta\mathcal{H}:=\{h(x)\oplus h^{\prime}(x)|h,h^{\prime}\in\mathcal{H}\}$ given $\oplus$ as a xor operation (proof is in Appendix B). We see that performance gap between two lanauges is bounded from above with a distribution divergence plus an irreducible term defined by problem setup. That is, if we reduce the divergence between language distributions, the expected performance gap can also be reduced accordingly.
To investigate whether this is indeed a case or not, we provided embedding space similarity and logit-space Sinkhorn distance between different languages in Figure 3. We argue that phonemic representation’s relatively mild performance gap is achieved by reducing linguistic gaps which is confirmed in the embedding space (high CKA) and final output space (low Sinkhorn distance).
5 Conclusion
Towards robust multilingual language modeling, we argue that mitigating the linguistic gap between different languages is crucial. Moreover, we advocate the use of IPA phonetic symbols as a universal language representation that partially bridges such linguistic gaps without any complicated cross-lingual training phase. Empirical validation on three representative NLP tasks demonstrates the superiority of phonemic representation compared to subword and character-based language representation in terms of the cross-lingual performance gap and linguistic gap. Theoretical analysis of the cross-lingual performance gap explains such promising results of phonemic representation.
6 Limitations
While we have shown that phonemic representation induces a small cross-lingual linguistic gap, therefore a small performance gap, the absolute performance of this phonemic representation is still lacking compared to subword-level models. We spur the necessity of putting research attention to developing phoneme-based LMs. Moreover, there is no such large phonemic language model beyond the BERT-base-size architecture, so we confine the scope of our empirical validation to BERT-base-size LMs. Thus, we can not ensure the effectiveness of IPA representation when adopted within modern large language models, such as LLaMa 2 Touvron et al. (2023). Additionally, we performed evaluation with a limited languages (up to 8), so it is unclear whether IPA language representations are effective for other numerous languages (especially low-resource ones) or not.
7 Ethics Statement
We believe there are no potential of any critical issues that harm the code of ethics provided by ACL. The social impacts of the technology—reducing performance gaps for low resource languages—will be, on the balance, positive. The data was, to the extent we can determine, collected in accordance with legal and institutional protocals.
Appendix A Details on Experiment Setup
A.1 Models
We compare LMs with three different types of language representation: subword, character, and phoneme. All models555pre-trained weights were obtained from https://huggingface.co/models are pre-trained on multilingual data that covers around 100 languages from Wikipedia dump files. We employ an off-the-shelf MLLMs, mBERT Devlin et al. (2018). for a subword-based model which is trained on 104 languages with byte-pair encoding (vocabulary size of 110k). The model is pre-trained with the objective of masked language modeling and next sentence prediction tasks. For a character-based model, we utilize CANINE Clark et al. (2022), which is a tokenization-free LM that directly maps each character to its codepoint by hashing. It is pre-trained on the same data and training objectives as mBERT. This prevents unknown tokens, enabling the model to handle a large amount of distinct characters. CANINE is pre-trained on the same data and training objectives as mBERT. For a phoneme-based model, we adopt XPhoneBERT Nguyen et al. (2023), which has the same model architecture as mBERT and it is trained on 94 languages with learning objective of dynamic masked language modeling.
While character-level models are known to better generalize to low-resource languages Clark et al. (2022), their general performance falls behind subword-based models. For a fair comparison between the representations, we compare phoneme-based model to character-based model instead of directly comparing it to widely used subword-based model (i.e., mBERT), and leave further improvements of overall performance as future work.
A.2 Data
G2P Conversion.
In order to pass the input to phoneme-based model, we employed G2P (Grapheme-to-Phoneme) conversion on the data to obtain an IPA version of the input. This conversion was done with Epitran [^6] https://github.com/dmort27/epitran Mortensen et al. (2018) , an external tool for G2P conversion. After converting to IPA, we insert white space between every character to make it compatible with the XPhoneBERT tokenizer.
Languages.
We categorize languages into three groups—high-resource (eng, deu, fra), low-resource (urd, hin, swa), and typologically distinct (kor, ukr)—to analyze the impact of phonemic representation with respect to each category. For high/low-resource language, we refer to Wu and Dredze (2020) and treat the languages with a wikisize under 8 as low-resource languages, and those with wikisize above 11 high-resource languages. On the other hand, typologically distinct languages are chosen with reference to English (they are orthographically and typologically different) [^7] Note that this categorization is not mutually exclusive and that Urdu and Hindi can also be considered as typologically distant from English..
A.3 Downstream Tasks
We adopt the cross-lingual generalization benchmark tasks suggested in XTREME Hu et al. (2020). Among them, we selected two types of tasks; classification and structured prediction.
Sentence-level Classification.
XTREME supports some sentence-level classification tasks, requiring semantic understanding of given sentences to make a prediction. We employ the XNLI Conneau et al. (2018) corpus, which is well-known for cross-lingual evaluation, to train and evaluate the model on different languages.
Token-level Classification.
Structured prediction tasks from Hu et al. (2020) include POS tagging and NER. Both tasks require labeling each token from the model. These types of tasks were previously analyzed as being relatively independent from the data size of each language used for pre-training Hu et al. (2020). We find this particularly suitable in our scenario where two models with different pre-training strategy are compared. We utilize the corpora from Universal Dependencies [^8] https://universaldependencies.org/ , v2.13, 148 languages, released Nov 15, 2023. for POS tagging, and WikiAnn Pan et al. (2017) with train, dev, test splits following Rahimi et al. (2019) for NER. While all 8 languages are employed for NER, we do not consider swa for POS tagging since Universal Dependencies does not provide a Swahili treebank.
Evaluation.
For evaluation, we follow the common practice of each task, evaluating XNLI task with top 1 accuracy and NER, POS tagging with F1 score. All metrics used are computed via functions from scikit-learn  [^9] https://scikit-learn.org/stable/index.html python library.
A.4 Implementation details.
We fine-tuned all models with AdamW Loshchilov and Hutter (2018) optimization setting batch size as 128 over 20 epochs for XNLI and 40 epochs for NER and POS Tagging tasks. Here, we adopt a cosine learning rate scheduler with a warm-up on the XNLI task. For all models, we adopt mixed precision training Micikevicius et al. (2018) provided by PyTorch, i.e., with autocast() loop for saving computational cost and energy consumption. For XNLI task, we used the embedding of [CLS] token for mBERT, and used the mean of all tokens except [CLS] token for CANINE and XPhoneBERT. All experiments were done with the random seed fixed to 42.
A.5 Number of parameters of the models.
Model Num. of params
mBERT-base 177,853,440
CANINE-c 132,082,944
XPhoneBERT 87,554,304
Table 3: The number of parameters of each model.
A.6 Computational resources.
All experiments on POS tagging were done on a single NVIDIA GeForce RTX 3090, and those on XNLI and NER were done on a single NVIDIA A6000. Total GPU hours for all experiments are 29 days.
A.7 Hyperparameter sweep.
We sweep hyperparameters over grid below (in Table 4), and select the final parameters for each model based on the best validation performance (Accuracy for XNLI and F1-score for NER and POS Tagging).
Task Hyperparam Search space Selected parameter value
mBERT CANINE XPhoneBERT
XNLI learning rate [5e-6, 7e-6, 1e-5, 3e-5, 5e-5] 5e-6 5e-6 (EN), 1e-5 (SW, UR) 7e-6 (EN), 3e-6 (SW, UR)
weight decay [0.0, 1e-1, 1e-2, 1e-3] 0.01 0.1 (EN), 0.0 (SW), 0.01 (UR) 0.1 (EN), 0.0 (SW), 0.01 (UR)
learning rate scheduling [True, False] True True False
NER learning rate [1e-5, 5e-5, 1e-4] 1e-4 5e-5 5e-5
weight decay [1e-4, 1e-3, 1e-2] 1e-3 1e-4 1e-2
POS learning rate [1e-5, 5e-5, 1e-4, 3e-4, 1e-2] 1e-4 3e-4 1e-4
weight decay [1e-4, 1e-3, 1e-2] 1e-4 1e-2 1e-2
Table 4: List of hyperparameter, search spaces and selected parameter values for different models applied to XNLI, NER, and POS tasks, detailing learning rate, weight decay, and learning rate scheduling for mBERT, CANINE, and XPhonemBERT, with specific configurations for optimal model performance per task.
A.8 Datasets, Statistics, and License.
In Table 5, we provide the datasets, their statistics, and license.
Dataset Lang. Train Dev Test License
XNLI eng 393k 2.49k 5.01k CC BY-NC-4.0
swa
urd
WikiAnn eng 20k 10k 10k ODC-BY
deu 20k 10k 10k
fra 20k 10k 10k
kor 20k 10k 10k
ukr 20k 10k 10k
urd 20k 1k 1k
swa 1k 1k 1k
hin 5k 1k 1k
UD eng 12.5k 2k 2k CC BY-SA 4.0
deu 13.8k 0.8k 1k CC BY-SA 4.0
fra 14.5k 1.5k 0.4k CC BY-SA 4.0
kor 23k 2k 2.3k CC BY-SA 4.0
ukr 5.5k 0.7k 0.9k CC BY-NC-SA 4.0
urd 4k 0.6k 0.5k CC BY-NC-SA 4.0
Table 5: Statistics and license types for datasets. The table lists the number of examples in the training, development, and testing sets for languages in the XNLI, WikiAnn, and UD datasets. It specifies the licensing conditions: CC-BY permits sharing and adaptation, CC BY-NC is for non-commercial usage, CC BY-SA mandates share-alike distributions, and CC BY-NC-SA combines non-commercial with share-alike terms. All datasets are strictly used within the bounds of these licenses.
Appendix B Details on Theoreoretical Analysis
We aim to diminish the performance gap between different languages by adopting IPA as a universal language representation. Motivated by domain adaptation literature Kifer et al. (2004); Ben-David et al. (2010), we present a theoretical justification of IPA for robust multilingual modeling by providing a bound for cross-lingual performance gap.
Let $\mathcal{D}$ denote a domain as a distribution over text feature input $\mathcal{X}$, such as the sequence of word embeddings or one-hot vectors, and a labeling function $f:\mathcal{X}\rightarrow\{0,1\}$. Assuming a binary classification task, our goal is to learn a hypothesis $h:\mathcal{X}\rightarrow\{0,1\}$ that is expected to minimize a risk $\varepsilon_{D}(h,f):=\mathbb{E}_{x\sim\mathcal{D}}[\mathbb{I}(f(x)\neq h(x))]$ and has a small risk-deviation over two domains $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$. Then, to formalize the cross-lingual performance gap, we first need a discrepancy measure between two languages. By following Ben-David et al. (2010), we adopt $\mathcal{H}$-divergence to quantify the distance between two language distributions.
Definition B.1 ($\mathcal{H}$-divergence; Ben-David et al. (2006)).
Let $\mathcal{H}$ be a hypothesis class for input space $\mathcal{X}$ and a collection of subsets from $\mathcal{X}$ is denoted by $\mathcal{S}_{\mathcal{H}}:=\{h^{-1}(1)|h\in\mathcal{H}\}$ which is the support of hypothesis $h\in\mathcal{H}$. The $\mathcal{H}$-divergence between two distributions $\mathcal{D}$ and $\mathcal{D^{\prime}}$ is defined as
$d_{\mathcal{H}}(D,D^{\prime})=2\sup_{S\in\mathcal{S}_{\mathcal{H}}}|\mathbb{P}_{D}(S)-\mathbb{P}_{D^{\prime}}(S)|$
$\mathcal{H}$-divergence is a relaxation of total variation between two distributions, and it can be estimated by finite samples from both distributions if $\mathcal{H}$ governs a finite VC dimension. Now, based on Lemma 1 and 3 of Ben-David et al. (2010), we make reasoning on performance gap over different language domains.
Theorem B.2.
Let $h:\mathcal{X}\rightarrow[0,1]$ be a real-valued function in a hypothesis class $\mathcal{H}$ with a pseudo dimension $\mathcal{P}dim(\mathcal{H})=d$. If $\hat{\mathcal{D}}_{A}$ and $\hat{\mathcal{D}}_{B}$ are the empirical distribution constructed by $n$-size i.i.d. samples, drawn from $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ respectively, then for any $\delta\in(0,1)$, and for all $h$, the bound below hold with probability at least $1-\delta$.
$|\varepsilon_{\mathcal{D}_{A}}(h,f)-\varepsilon_{\mathcal{D}_{B}}(h,f)|\leq{1\over 2}d_{\mathcal{H}\Delta\mathcal{H}}(\hat{\mathcal{D}}_{A},\hat{\mathcal{D}}_{B})$
$+2\sqrt{{{d\log(2n)+\log(2/\delta)}\over n}}$
where $\mathcal{H}\Delta\mathcal{H}:=\{h(x)\oplus h^{\prime}(x)|h,h^{\prime}\in\mathcal{H}\}$ given $\oplus$ as a xor operation.
proof of Theorem B.2..
we start to prove Theorem B.2. by restating Lemma 1 of Ben-David et al. (2010) adapted to our notation.
Lemma B.3.
Let $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ be distributions of domain $A$ and $B$ over $\mathcal{X}$, respectively. Let $\mathcal{H}$ be a hypothesis class of functions from $\mathcal{X}$ to $[0,1]$ with VC dimension d. If $\hat{\mathcal{D}}_{A}$ and $\hat{\mathcal{D}}_{B}$ are the n-size empirical distributions generated by $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ respectively, then, for $10<\delta<$ 1, with probability at least $1-\delta$,
$d_{\mathcal{H}}(\mathcal{D}_{A},\mathcal{D}_{B})$ $\leq d_{\mathcal{H}}(\hat{\mathcal{D}}_{A},\hat{\mathcal{D}}_{B})$
$+4\sqrt{{{d\log(2n)+\log(2/\delta)}\over n}}.$
Then, for any hypothesis function $h,h^{\prime}\in\mathcal{H}$, by the definition of $d_{\mathcal{H}\Delta\mathcal{H}}$-divergence, we have:
$d_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{D}_{A},\mathcal{D}_{B})$
$=2\sup_{h,h^{\prime}\in\mathcal{H}}|\mathbb{P}_{x\sim\mathcal{D}_{A}}[h(x)\neq h^{\prime}(x)]-\mathbb{P}_{x\sim\mathcal{D}_{B}}[h(x)\neq h^{\prime}(x)]|$
$=2\sup_{h,h^{\prime}\in\mathcal{H}}|\varepsilon_{\mathcal{D}_{A}}(h,h^{\prime})-\varepsilon_{\mathcal{D}_{B}}(h,h^{\prime})|$
$\geq 2|\varepsilon_{\mathcal{D}_{A}}(h,h^{\prime})-\varepsilon_{\mathcal{D}_{B}}(h,h^{\prime})|$
Now the below bound holds for any hypothesis functions $h,h^{\prime}\in\mathcal{H}$ (See Lemma 3 of Ben-David et al. (2010)).
$|\varepsilon_{\mathcal{D}_{A}}(h,h^{\prime})-\varepsilon_{\mathcal{D}_{B}}(h,h^{\prime})|\leq{1\over 2}d_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{D}_{A},\mathcal{D}_{B})$
Finally, by plugging the Lemma B.3 into the above bound, we have Theorem B.2.
From Theorem B.2, we see that the difference between true risks across language domains is bounded by an empirical estimation of the divergence ($d_{\mathcal{H}\Delta\mathcal{H}}$) between those two domains plus an irreducible term defined by problem setup. Thus, if we reduce the divergence between language distributions, the expected performance gap can also be reduced accordingly. To investigate whether this is indeed a case or not, we provided the embedding-space similarity and logit-space Sinkhorn distance between different languages in Figure 3. We argue that phonemic representation’s relatively mild performance gap is achieved by reducing linguistic gaps in the embedding space (high CKA) and final output space (low Sinkhorn distance) those are the proxy of $\mathcal{H}$-divergence.